OpenStudy (anonymous):

A student has the following grades: 86, 80, 79, 85, 70. Find the standard deviation of the grades. 1) 1640.5 2) 5.69 3) 40.5 4) 32.4

OpenStudy (amistre64):

sd = sqrt(variance)

OpenStudy (amistre64):

\[\sigma = \sqrt{\frac{\sum (x-\mu)^2}{n}}\]

OpenStudy (anonymous):

@tana Should you just enter into the calculator, use the Stat - var char option!

OpenStudy (anonymous):

its 32.4

OpenStudy (anonymous):

I don't have the appropriate calculator now!

OpenStudy (amistre64):

its the sqrt(32.4)